OpenStudy (anonymous):

A 27 story skyscrapper is 550 feet high with 400,000 square feet of living space. What would be the cost of carpeting the entire living space if the cost of the carpet were $36 per square yard?

OpenStudy (campbell_st):

well how many square feet in a square yard..?

OpenStudy (anonymous):

one yard is 3 feet so one square yard is \(3^2=9\) square feet divide \(400,000\) by 9, then multiply the result by \(36\) to get the cost

OpenStudy (anonymous):

in other words, compute \[\frac{36\times 400,000}{9}=4\times 400,000\]